40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 12 00:47:45 2013, Den, Update, LSC, locking activity POP_ASC.pdf
    Reply  Fri Dec 13 02:19:57 2013, Jenne, Update, LSC, locking activity PRMI_Yarm_NewPRCLasc.pngtry_dc.pdf
       Reply  Fri Dec 13 02:30:22 2013, Koji, Update, LSC, locking activity 
Message ID: 9462     Entry time: Fri Dec 13 02:19:57 2013     In reply to: 9456     Reply to this: 9463
Author: Jenne 
Type: Update 
Category: LSC 
Subject: locking activity 

[Jenne, Den]

Tonight we worked on tweaking up the PRCL new ASC, and then PRMI+1 arm locking.  We were unable to get the Xarm to stay locked on a TEM00 mode for very long, and after an hour or two of using the PZTs to try to align the beam to the cavity, we gave up and just used Yarm green.

NB: We haven't done anything to MCL, although it is not in use.  Den is still going to get around to elogging what servo shaping he changed on that last night.

I wrote a script that will handle the transitions between the new PRCL ASC and the PRM oplev and local damping.  The script is accessible from the PRC ASC screen, and will detect when the PRMI is locked or not.  When it is locked, it will turn down the PRM oplev gains and turn on the ASC, and then it will turn off the local shadow sensor damping for PRM pitch and yaw.  When the PRMI unlocks, the script will turn off the ASC and restore oplev and local shadow sensor damping. 

We saw that the bounce mode of the PRM was getting rung up with our new ASC, so we included a band stop in the ASC, and also turned on the triggering for the PRCL LSC FM6, which has the resonant gain for the bounce mode (as well as roll, and the stack mode).  This made the PRMI spot very stable. 

We then moved on to green arm locking.  The Yarm is behaving perfectly nicely (as nice as it has been lately - it's alignment and mode matching could also use some work), but Xarm was giving us a bit of trouble.  As always (since the PZTs were installed?), the mode matching isn't excellent for the green to the arm, so it can be hard to catch a TEM00 mode.  Also, even if we did catch a good mode, it would often not stay locked for more than a few tens of seconds.  We tried several alignment tweakings, and several different end laser temperatures (within the confines of seeing the beatnote under 100MHz), and didn't have a lot of success.  It looks like Eric had the slow servo engaged for the Xend laser, so the temperature offset was something like +300,000, which seemed totally crazy.  I turned that off, and found the beatnote somewhere around output of -10,300.  So, I haven't gone to the end to look at the temperature, but it's going to be different than when Eric was taking measurements this afternoon.  It seems like the main problem with the Xarm is poor mode matching - the maximized input pointing for TEM00, when you unlock and relock the cavity, is just as likely to give you a TEM_9_0 mode, as TEM00.

So, we gave up on the Xarm for the evening, and tried PRMI+1arm, with the new PRCL ASC.  This was successful!   The Yarm beatnote was around laser slow servo output of +4450.  Beatnote at 46.0MHz, -26dBm.  We found the IR resonance, moved off, locked the PRMI, transitioned to the new ASC, and brought the Yarm back to IR resonance.  What we see is that the power fluctuations in the PRC are much smaller than they were back around Halloween (elog 9338), however the arm power fluctuations still seem very, very large.  This is certainly partly due to the fact that we haven't done a thorough Yarm alignment since before messing with the greens, so we will have drifted somewhat.  Also, the ALS beatnote sensor isn't perfect, so won't be perfect at holding us near resonance. 

Den is thinking about whether we can use the arm transmission QPD signals to feed back to the ETM ASC servos, to try to reduce the RIN in the arms.  I feel like we should also see if this amount of power fluctuation can be explained by our ALS noise, because maybe we'll be fine once we transition to IR and turn off the ALS system.   Attached is a plot showing that the arm's RIN is coherent with the spot motion seen by the transmission QPD, so we need to check the alignment of the cavity, as well as consider using the trans QPD in an ASC feedback loop.

Here is a plot of the PRC sideband power, as well as the Yarm transmission.  The GPS time for this plot is approximately 1070963372.

PRMI_Yarm_NewPRCLasc.png

Attachment 2: try_dc.pdf  36 kB  | Hide | Hide all
try_dc.pdf
ELOG V3.1.3-